OpenStudy (anonymous):

use a half- angle formula to find the exactly value of the expression...: Sin5 pies/12... can someone help me step by step

OpenStudy (anonymous):

Do you know the half-angle formulas?

OpenStudy (anonymous):

yes, i do...

OpenStudy (anonymous):

What's the formula for sin(x/2)?

OpenStudy (anonymous):

sin(x/2)=\[\pm \sqrt{1-cosx/2}\]

OpenStudy (anonymous):

Ok. You missed the parentheses around 1-cosx, but yes that checks out. Now, you know how to find the sin of half of an angle. You need to find sin(5pi/12). Can you see how this formula might help?

OpenStudy (anonymous):

sin(x/2)=\[\pm \sqrt{1-cosx/2}\]

OpenStudy (anonymous):

i dont know how to figure out:)..

OpenStudy (anonymous):

sin(x/2)=\[\pm \sqrt{1-cosx/2}\]

OpenStudy (anonymous):

in the unit circle dont have 5pi/12

OpenStudy (anonymous):

They give you a hint to use the half-angle formula. What is 5pi/12 half of?

OpenStudy (anonymous):

75 degree right

OpenStudy (anonymous):

Yes, but leave it in terms of pi.

OpenStudy (anonymous):

Wait. 75 degrees is still half of the angle we want. And keep it in radians. So 5pi/12 is half the angle we want. So what angle is 5pi/12 half of?

OpenStudy (anonymous):

5pi/24

OpenStudy (anonymous):

Wrong way.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

5pi/12 is half of the angle I've marked with a ?. How big is that angle?